OpenStudy (anonymous):

How do u find sin(arccos1/2+arcsin-1)

OpenStudy (campbell_st):

well arccos(1/2) = 60 or pi/3 arcsin(-1) = 270 or 3pi/2 so you are looking at sin(60 + 270) = sin(330) = -sin(30)

OpenStudy (campbell_st):

just write it as an exact value

OpenStudy (anonymous):

Thank you so much(: